# Categories of Line Defects and Cohomological Hall Algebras

### v1

@@ 5.1 Framed quivers and BPS degeneracies
We are led to conjecture the existence of a monoidal functor $\mathrm{RG}$ mapping $L_{m,e}$ to the bimodule with $\gamma_{*}=(m,e)$ and thus $f_{1}=m$. In order to test this conjecture, we compute tensor products of $B_{\pm 1}$, corresponding to $\gamma_{*}=(\pm 1,0)$.
@@
We conclude that the functor $\mathrm{RG}$ remembers the extension in $L_{1,0}\cdot L_{-1,0}$ but not the extension in $L_{-1,0}\cdot L_{1,0}$.
We could repeat the analysis in the second chamber, but it is easier to employ the $\mathbb{Z}_{2}$ symmetry (159), which permutes the two chambers. Then we see that in the second chamber, the functor $\mathrm{RG}$ remembers the extension in $L_{-1,0}\cdot L_{1,0}$ but not the extension in $L_{1,0}\cdot L_{-1,0}$.
In conclusion, the monoidal structure we found is compatible with the existence of two monoidal functors $\mathrm{Lines}\to\mathrm{sH}(A_{1})-\mathrm{Bimod}$, one per chamber: each functor “detects” one of the two non-trivial extensions and forgets the other. Together, they detect the whole extension structure.181818 Amusingly, this match has the shape of a Koszul duality: the $\mathrm{Lines}$ calculations roughly employ interfaces in a B-model on the product of infinitely many copies of $\mathbb{C}$, which can be described as coherent sheaves but also, via Koszul duality, as bimodules for an infinite-dimensional fermionic algebra. It would be interesting to develop this point further.
@@ 6.1 $\mathbb{Z}_{5}$ Orbits of Framed Quivers
$F(B_{3})$ $=X_{-\gamma_{2}}+X_{\gamma_{1}}+X_{\gamma_{1}-\gamma_{2}}\,,$ (202)
@@ 7.2 Ring relations
and extend the labeling by a modified $\mathbb{Z}_{3}$ periodicity
@@ Appendix A Line defects as modules of quiver Yangian
One of the main results of the current paper is the description of the line defects as the CoHA-binomule from the corresponding framed quiver, see Section. 4. In this appendix, we compare this with an alternative description of the line defects as modules of the so-called quiver Yangian, in terms of the ideals of the Jacobian algebra of the framed quiver Li:2023zub .
@@ A.2 Higher order relations from quiver Yangian arguments
A.2.1 Three-loop quiver
@@
Let’s now apply this argument on the three-loop quiver. The generators are
@@ B.1 SUSY traps
We will now deform the problem in a manner which “traps” the particle at some location and thus gives a better notion of SUSY ground states and $Q$-cohomology. We are looking for deformations which are as “canonical” as possible, in the sense that they arise from some universal type of deformation in the original 4D theory.

### v2

@@ 5.1 Framed quivers and BPS degeneracies
We are led to conjecture the existence of a monoidal functor $\mathrm{RG}$ mapping $L_{m,e}$ to the bimodule with $\gamma_{*}=(m,e)$ and thus $f_{1}=m$. In order to test this conjecture, we compute tensor products of $B_{\pm 1}$, corresponding to $\gamma_{*}=(\pm 1,0)$. We will first ignore the $\Gamma^{\vee}({\mathcal{O}})$ equivariance and then amend that.
@@
At thus point it would appear that the functor $\mathrm{RG}$ remembers the extension in $L_{1,0}\cdot L_{-1,0}$ but not the extension in $L_{-1,0}\cdot L_{1,0}$.
Luckily, the splitting of $B_{-1}\otimes_{A}B_{1}$ is not compatible with the action of $\Gamma^{\vee}({\mathcal{O}})={\mathcal{O}}^{*}$ on the bimodules. Recall that the non-trivial generators of ${\mathcal{O}}^{*}$ act as
$o_{k}:\,e_{n}\to e_{n}+e_{n+k}\,,$ (184)
see (55). Accordingly,
$o_{k}:\,b_{-1}\otimes e_{0}b_{1}\to b_{-1}\otimes e_{0}b_{1}+b_{-1}\otimes b_{1}e_{k-1}$ (185)
This is an extension! In other words, $A$ is an ${\mathcal{O}}^{*}$-equivariant submodule of the tensor product, but $\mathfrak{q}^{-1}X_{\gamma_{1}}A$ is not and we get a non-trivial exact sequence
$0\to A\to B_{-1}\otimes_{A}B_{1}\to\mathfrak{q}^{-1}X_{\gamma_{1}}A\to 0\,.$ (186)
It appears that keeping track of ${\mathcal{O}}^{*}$ equivariance allows for an exact match of the monoidal structure in $\mathrm{Lines}$!
We could repeat the analysis in the second chamber, but it is easier to employ the $\mathbb{Z}_{2}$ symmetry (170), which permutes the two chambers.
In conclusion, the monoidal structure we found is compatible with the existence of two monoidal functors
$\mathrm{Lines}\to\mathrm{sH}(A_{1})-\mathrm{Bimod}_{{\mathcal{O}}^{*}}\,,$ (187)
one per chamber, which detect the extension structure of the tensor products of $L_{1}$ and $L_{-1}$.
As $\mathrm{Lines}=\mathrm{KP}_{\mathbb{C}^{*},\mathbb{C}}$, it should be possible to verify the existence of the conjectured quasi-isomorphism $\mathrm{RG}_{u}$. We will do so in upcoming work.
@@ 6.1 $\mathbb{Z}_{5}$ Orbits of Framed Quivers
$F(B_{3})$ $=X_{-\gamma_{2}}+X_{\gamma_{1}-\gamma_{2}}+X_{\gamma_{1}}\,,$ (217)
@@ 6.3 Interesting tensor products
All the maps we describe are $\Gamma^{\vee}({\mathcal{O}})$-equivariant (sometimes non-trivially).
@@ 7.2 Ring relations
and extend the labelling by a modified $\mathbb{Z}_{3}$ periodicity
@@ Appendix A Line defects as modules of quiver Yangian
One of the main results of the current paper is the description of the line defects as the CoHA-bimodule from the corresponding framed quiver, see Section. 4. In this appendix, we compare this with an alternative description of the line defects as modules of the so-called quiver Yangian, in terms of the ideals of the Jacobian algebra of the framed quiver Li:2023zub .
@@ A.2 Higher order relations from quiver Yangian arguments
A.2.1 Three-node-one-loop quiver
@@
Let’s now apply this argument on the three-node-one-loop quiver. The generators are
@@ B.1 SUSY traps
We will now deform the problem in a manner which “traps” the particle at some location and thus gives a better notion of SUSY ground states and $Q$-cohomology. [^28] For related discussion, see Cecotti:2011iy We are looking for deformations which are as “canonical” as possible, in the sense that they arise from some universal type of deformation in the original 4D theory.